# UniTS: A Unified Multi-Task Time Series Model

### v2

@@ 6 Conclusion
We have developed UniTS, a unified model for time series analysis that employs a universal specification of time series tasks. It effectively handles multi-domain time series data with heterogeneous representations, outperforming task-specific models and reprogrammed LLMs on 38 multi-domain and multi-task datasets. Additionally, UniTS demonstrates strong zero-shot, few-shot, and prompt-based performance, indicating its ability to generalize to new domains and tasks. Limitations and future directions are discussed in Appendix L.
@@ Acknowledgments
S.G., O.Q., and M.Z. gratefully acknowledge the support of NIH R01-HD108794, NSF CAREER 2339524, US DoD FA8702-15-D-0001, awards from Harvard Data Science Initiative, Amazon Faculty Research, Google Research Scholar Program, AstraZeneca Research, Roche Alliance with Distinguished Scientists, Sanofi iDEA-iTECH Award, Pfizer Research, Chan Zuckerberg Initiative, John and Virginia Kaneb Fellowship award at Harvard Medical School, Aligning Science Across Parkinson’s (ASAP) Initiative, Biswas Computational Biology Initiative in partnership with the Milken Institute, and Kempner Institute for the Study of Natural and Artificial Intelligence at Harvard University. They also acknowledge the significant contributions of the community in providing open-source datasets.
@@ Appendix B Datasets
Dataset details. We introduce the details of the multi-task dataset collection used by our work in Table 7. The dataset collection used for few-shot learning on classification and forecasting are listed in Table 8, the collection used for zero-shot forecasting are listed in Table 9, the collection used for imputation is listed in Table 10, and the collection used for anomaly detection is listed in Table 11. Datasets were aggregated from the Monash Forecasting Repository [33], Time Series Classification Website [77], and Time Series Library [109]. The combined training set consists of over 35 million timesteps and over 6,000 variables.
Dataset for zero-shot learning on new forecasting lengths. For evaluating zero-shot learning capabilities over new forecasting lengths, we initially consider 20 forecasting datasets utilized in the multi-task setting, as detailed in Table 7. However, to adapt to 384 additional forecasting lengths that the model was not trained on, we exclude specific datasets that are incompatible with this requirement. These datasets include NN5P112, ECLP720, ETTh1P720, ILIP60, TrafficP720, and WeatherP720. Consequently, our analysis is conducted using 14 remaining forecasting datasets.
@@ D.2 Training Details
For multi-task settings, all models are jointly trained on multiple tasks following the same training protocol. To match the size of the largest dataset, samples from each dataset are repeated in every training epoch. In each inference step, datasets are randomly sampled with equal probability, utilizing a batch size of 32. Supervised training involves 5 epochs using gradient accumulation for an effective batch size of 1024, starting with a learning rate of 3.2e-2 and adjusted with a multi-step decayed schedule. The $\lambda_{i}$ in $L_{\text{total}}$ are all set to 1 in this work. For self-supervised pretraining, the models are trained over 10 epochs with an effective batch size of 4096 and an initial learning rate of 6.4e-3, using a cosine decay schedule. All experiments are conducted using A100-40G GPUs. Each experiment is conducted with one or two GPUs, and the maximum running time is under 48 hours.
@@
D.3 Further Information on Pretraining
During the unified pretraining, we introduce two distinct masking schemes: the random masking scheme and the right masking scheme. The time series sample is initially truncated to a length randomly selected within the range of 50% to 100% of its original length. Subsequently, in the random masking scheme, a certain proportion $p_{\text{rand}}$ of tokens are masked at random positions within the time dimension. For the right masking scheme, designed to enhance the model’s forecasting ability, a random proportion $p_{\text{right}}$ of tokens on the right side of the sample is masked. Both $p_{\text{rand}}$ and $p_{\text{right}}$ are set to 70%-80%. Each training step randomly utilizes one of these two schemes with equal probability.
@@ D.4 Implementation Details of Baselines
The baseline methods used in this paper are summarized in Table 13. For multi-task learning settings, models are configured with 3 blocks, except for GPT4TS, which utilizes the prescribed setting of 6 GPT blocks. For both the proposed method and patch-based baseline approaches, the patch size and stride are fixed at 16. The input and output heads of baseline methods are duplicated for each task to create data/task-specific heads tailored for each data source and task. For single-task learning settings, we follow the original settings of baseline methods and compare results reported in their papers.
Appendix E Additional Results: Prompt Learning and Pretraining
We do more analysis on the prompting and pretraining of UniTS. The average performance under 38 datasets with the multi-task setting is reported.
@@ Appendix E Additional Results: Prompt Learning and Pretraining
Table 17: Ablation on the pretraining scheme.
@@
Unified pretraining. In Equation 5, the proposed unified mask reconstruction pretraining loss is detailed, consisting of two components: the mask reconstruction loss associated with prompt tokens and the mask reconstruction loss related to CLS tokens. Table 17 presents the results where either the CLS token-based reconstruction loss or the prompt token-based reconstruction loss is omitted. The performance of prompt learning is reported. The results highlight the impact of each loss component on the learning performance.
Specifically, excluding the CLS token-based loss resulted in a significant decline in classification performance, dropping sharply from 78.0% to 33.1%. This substantial drop underscores the critical role of the CLS token-based pretraining loss in enabling the model’s classification capabilities. Conversely, the removal of the prompt token-based loss adversely affected the forecasting performance. For instance, the MSE drops from 0.471 to 0.967. This deterioration in performance demonstrates the importance of prompt token-based pretraining in generative tasks.
@@ Appendix F Additional Results: Ablation Studies of UniTS
Table 19: Ablation on the Dynamic FFN layer in UniTS network.
@@
Effect of Dynamic FFN. In Table 19, we present an ablation study on the Dynamic FFN layer in the UniTS network. The UniTS, which incorporates the Dynamic FFN, achieves the highest performance with an average accuracy of 81.6%, demonstrating effectiveness in handling classification tasks. It also shows superior results in terms of MSE and MAE in forecasting tasks, with scores of 0.439 and 0.381 respectively. The model variant where the Dynamic FFN is replaced with a standard FFN layer exhibits a decrease in performance. The average accuracy dropped to 81.3%, and MSE and MAE dropped to 0.462 and 0.394, respectively. This variation suggests the effect of Dynamic FFN for the UniTS. The performance is observed when the Dynamic FFN is completely removed from the model, highlighting the importance of Dynamic FFN layers in UniTS network.
@@ Appendix G Additional Results: UniTS for Zero-Shot Forecasting on New Datasets
Setup. When UniTS is trained with shared prompt and GEN tokens across all forecasting tasks, it acquires the ability to perform zero-shot forecasting on datasets with new lengths and variable numbers that were not part of its training domain. We evaluate UniTS in a zero-shot setting on five new forecasting tasks as referenced in Table 9. These tasks have varying forecasting lengths and numbers of variables compared to those seen by UniTS during pretraining. We benchmark against LLMTime [79], a model designed for zero-shot forecasting using LLMs. Following LLMTime, we utilize one sample from each dataset to manage the extensive inference costs. We exclude a related method, Time-LLM [46], from experiments. Time-LLM supports zero-shot learning but requires that the forecasting length and the number of variables/sensors for zero-shot prediction are the same as those used for training.
@@
Figure 6: The similarity of prompt tokens among datasets.
@@ Appendix I Additional Results: Classification Performance Stratified by Datasets
Appendix J Additional Results: Zero-Shot Learning on New Forecasting Lengths
@@
Appendix L Limitations and Future Directions
@@
The unified token scheme in UniTS allows it to represent data and tasks in a general manner. UniTS’s network employs a transformer-like architecture for simplification; and we plan to include other types of network structures, such MLP-based blocks [104, 14], to further enhance UniTS.
Appendix M Impact Statement

### v3

@@ 6 Conclusion
We have developed UniTS, a unified model for time series that uses a universal specification of time series tasks. UniTS handles multi-domain time series data with heterogeneous representations, outperforming task-specific models and reprogrammed LLMs on 38 multi-domain and multi-task datasets. UniTS also shows strong few-shot and prompt-based performance and can generalize to new domains and tasks. The unified token scheme in UniTS allows it to represent data and tasks in a general manner. UniTS uses a transformer architecture, and we plan to explore other types of backbones, such MLP-based blocks [107, 14] and Mamba [38], to further enhance UniTS. Limitations and future directions are discussed in Appendix M.
@@ Acknowledgments
S.G., O.Q., and M.Z. gratefully acknowledge the support of NIH R01-HD108794, NSF CAREER 2339524, US DoD FA8702-15-D-0001, awards from Harvard Data Science Initiative, Amazon Faculty Research, Google Research Scholar Program, AstraZeneca Research, Roche Alliance with Distinguished Scientists, Sanofi iDEA-iTECH, Pfizer Research, Chan Zuckerberg Initiative, John and Virginia Kaneb Fellowship at Harvard Medical School, Biswas Computational Biology Initiative in partnership with the Milken Institute, Harvard Medical School Dean’s Innovation Fund for the Use of Artificial Intelligence, and Kempner Institute for the Study of Natural and Artificial Intelligence at Harvard University. They also acknowledge the significant contributions of the community in providing open-source datasets. T.H. acknowledges the support of the National Security Data & Policy Institute, Contracting Activity 2024-24070100001. Any opinions, findings, conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the funders.
@@ Appendix B Datasets
Dataset details. We introduce the details of the multi-task dataset collection used by our work in Table 7. The dataset collection used for few-shot learning on classification and forecasting are listed in Table 8, the collection used for zero-shot forecasting are listed in Table 9, the collection used for imputation is listed in Table 10, and the collection used for anomaly detection is listed in Table 11. Datasets were aggregated from the Monash Forecasting Repository [33], Time Series Classification Website [79], and Time Series Library [112]. The combined training set consists of over 35 million timesteps and over 6,000 variables. For subsets of a dataset such as ETTh1, we start by splitting the data into training and testing sets based on distinct time intervals of a long time series sequence, following splits in [112]. Within these training and testing intervals, we generate samples using various sliding windows, ensuring that there is no data leakage between the training and testing sets.
Dataset for direct multi-step forecasting on new forecasting lengths. For evaluating zero-shot learning capabilities over new forecasting lengths, we initially consider 20 forecasting datasets utilized in the multi-task setting, as detailed in Table 7. However, to adapt to 384 additional forecasting lengths that the model was not trained on, we exclude specific datasets that are incompatible with this requirement. These datasets include NN5P112, ECLP720, ETTh1P720, ILIP60, TrafficP720, and WeatherP720. Consequently, our analysis is conducted using 14 remaining forecasting datasets.
@@ D.2 Training Details
For multi-task settings, all models are jointly trained on multiple tasks following the same training protocol. To match the size of the largest dataset, samples from each dataset are repeated in every training epoch. In each inference step, datasets are randomly sampled with equal probability, utilizing a batch size of 32. Supervised training involves 5 epochs using gradient accumulation for an effective batch size of 1024, starting with a learning rate of 3.2e-2 and adjusted with a multi-step decayed schedule. The $\lambda_{i}$ in $L_{\text{total}}$ are all set to 1 in this work. For self-supervised pre-training, the models are trained over 10 epochs with an effective batch size of 4096 and an initial learning rate of 6.4e-3, using a cosine decay schedule. All experiments are conducted using A100-40G GPUs. Each experiment is conducted with one or two GPUs, and the maximum running time is under 48 hours.
@@
D.3 Further Information on Pre-training
During the unified pre-training, we introduce two distinct masking schemes: the random masking scheme and the right masking scheme. The time series sample is initially truncated to a length randomly selected within the range of 50% to 100% of its original length. Subsequently, in the random masking scheme, a certain proportion $p_{\text{rand}}$ of tokens are masked at random positions within the time dimension. For the right masking scheme, designed to enhance the model’s forecasting ability, a random proportion $p_{\text{right}}$ of tokens on the right side of the sample is masked. Both $p_{\text{rand}}$ and $p_{\text{right}}$ are set to 70%-80%. Each training step randomly utilizes one of these two schemes with equal probability.
@@ D.4 Implementation Details of Baselines
The baseline methods used in this paper are summarized in Table 13. Unlike UniTS, which can handle diverse data and tasks within a single model, baseline methods cannot be directly used for unified training because: 1) To accommodate data with varying numbers of variables, baseline methods typically use a data-specific input head to project features from the variable count to a fixed number of embedding dimensions. 2) Similarly, to manage different tasks, such as classification with various classes and forecasting with different lengths, baseline methods employ task-specific output heads to transform the features into the appropriate task outputs. Since baseline methods are designed for single-task training, in their original setting, data/task-specific heads are used for each data and task. In the multi-task learning setting, to make baseline methods support unified training, we add separate input heads to project data into a shared embedding space and separate output heads to convert the shared model output into task-specific outputs. However, using separate input and output heads makes it hard to generalize to new datasets and tasks. We employ the same fully supervised multi-task training approach as UniTS. In this setting, model networks are stacked with 3 basic building blocks, except for GPT4TS, which utilizes the prescribed setting of 6 GPT blocks. For both the proposed method and patch-based baseline approaches, the patch size and stride are fixed at 16. The input and output heads of baseline methods are duplicated for each task to create data/task-specific heads tailored for each data source and task. For single-task learning settings, we follow the original settings of baseline methods and compare results reported in their papers.
Appendix E Additional Results: Prompt Learning and Pre-training
We do more analysis on the prompting and pre-training of UniTS. The average performance under 38 datasets with the multi-task setting is reported.
@@ Appendix E Additional Results: Prompt Learning and Pre-training
Table 17: Ablation on the pre-training scheme.
@@
Unified pre-training. In Equation 5, the proposed unified mask reconstruction pre-training loss is detailed, consisting of two components: the mask reconstruction loss associated with prompt tokens and the mask reconstruction loss related to CLS tokens. Table 17 presents the results where either the CLS token-based reconstruction loss or the prompt token-based reconstruction loss is omitted. The performance of prompt learning is reported. The results highlight the impact of each loss component on the learning performance.
Specifically, excluding the CLS token-based loss resulted in a significant decline in classification performance, dropping sharply from 78.0% to 33.1%. This substantial drop underscores the critical role of the CLS token-based pre-training loss in enabling the model’s classification capabilities. Conversely, the removal of the prompt token-based loss adversely affected the forecasting performance. For instance, the MSE drops from 0.471 to 0.967. This deterioration in performance demonstrates the importance of prompt token-based pre-training in generative tasks.
Pre-training with scaled numbers of epochs and data sizes. To evaluate the effect of scaling effect of pre-training, we conduct experiments of pre-training UniTS by varying the size of the pre-training dataset and the amount of training epochs. As demonstrated in Table 18, increasing the number of pre-training epochs improves performance on both forecasting and classification tasks. Similarly, increasing the size of pre-training dataset improves performance on both forecasting and classification tasks, as shown in Table 19.
Table 18: Performance of UniTS under different pre-training epochs, average performance on 20 forecasting and 18 classification are reported.
Pre-training steps 1 epoch 3 epochs 5 epochs 8 epochs 10 epochs
[…]
Table 19: Performance of UniTS under different pre-training data sizes, average performance on 20 forecasting and 18 classification are reported. Pre-training data size refers to the proportion of the total training set used.
Pre-training data size 10% 30% 50% 80% 100%
[…]
Cross-task pre-training. We evaluate the effect of cross-task pre-training by pre-training a model using our pre-training strategy on either generative tasks (forecasting) or predictive tasks (classification). Table 20 shows that UniTS, pre-trained solely on forecasting datasets, achieves similar performance to the model pre-trained on both forecasting and classification data. Despite not encountering any classification datasets during pre-training, it still performs well on classification tasks. When the model is pre-trained exclusively on classification datasets, performance on both classification and forecasting tasks drops significantly compared to the model pre-trained on both types of data. Given that the data amount of forecasting datasets is larger than classification datasets (22920 vs. 5022 iterations per epoch), this suggests that the larger amount of data plays a more crucial role in pre-training effectiveness than the data type.
Table 20: Cross-task pre-training evaluation on UniTS, average performance on 20 forecasting and 18 classification tasks are reported.
Evaluation data
[…]
20 forecasting datasets 78.5 0.454 0.379
18 classification datasets 74.1 0.583 0.807
Full 38 datasets 79.0 0.460 0.383
Cross-domain pre-training. We evaluate the effect of cross-domain data pre-training, where the model is pre-trained on either Weather-domain datasets or Traffic-domain datasets. In Table 21, compared to joint pre-training on both domains, the performance decreases with single-domain pre-training, where pre-training is conducted solely on the downstream dataset’s domain, showing the advantage of joint pre-training. For instance, the MSE on Weather datasets goes from 0.253 to 0.259. Compared to single-domain pre-training, cross-domain pre-training leads to larger performance drops, e.g., pre-training on Traffic datasets and then evaluating on Weather datasets results in an MSE increase from 0.259 to 0.289. Interestingly, pre-training on Weather datasets achieves better performance across both domains, suggesting that data from certain domains might be more beneficial for pre-training.
Table 21: Cross-domain pre-training evaluation on UniTS, average performance on 4 Weather or Traffic dataset domains are reported.
Weather datasets (4 sets) Traffic datasets (4 sets)
[…]
Weather domain (4 datasets) 0.259 / 0.287 1.338 / 0.768
Traffic domain (4 datasets) 0.289 / 0.314 0.680 / 0.438
Weather + Traffic domains (8 sets) 0.253 / 0.282 0.511 / 0.320
@@ Appendix F Additional Results: Ablation Studies of UniTS
Table 23: Ablation on the MLP layer in UniTS network.
@@
Effect of Dynamic FFN. In Table 23, we present an ablation study on the Dynamic FFN layer in the UniTS network. The UniTS, which incorporates the Dynamic FFN, achieves the highest performance with an average accuracy of 81.6%, demonstrating effectiveness in handling classification tasks. It also shows superior results in terms of MSE and MAE in forecasting tasks, with scores of 0.439 and 0.381 respectively. The model variant where the Dynamic FFN is replaced with a standard MLP layer exhibits a decrease in performance. The average accuracy dropped to 81.3%, and MSE and MAE dropped to 0.462 and 0.394, respectively. This variation suggests the effect of Dynamic FFN for the UniTS. The performance is observed when the Dynamic FFN is completely removed from the model, highlighting the importance of Dynamic FFN layers in UniTS network.
@@
Comparison with Transformer. To verify the effectiveness of UniTS structure, we compare the original Transformer with UniTS. The unified tokenization and co-training strategy are applied to both models. The results shown in Table 26 indicate that UniTS clearly outperforms the Transformer in both classification and forecasting tasks, suggesting that merely using a transformer structure is insufficient for achieving robust multi-task performance on time series datasets.
Table 26: Comparison between UniTS and Transformer structure. The unified tokenization and co-training strategy are applied to both models.
[…]
Transformer-network 80.2% 0.468 0.397
UniTS-network 81.6% 0.439 0.381
@@ Appendix G Additional Results: UniTS for Zero-Shot Forecasting on New Datasets
Setup. When UniTS is trained with shared prompt and GEN tokens across all forecasting tasks, it acquires the ability to perform zero-shot forecasting on datasets with new lengths and variable numbers that were not part of its training domain. We evaluate UniTS in a zero-shot setting on five new forecasting tasks as referenced in Table 9. These tasks have varying forecasting lengths and numbers of variables compared to those seen by UniTS during pre-training. We benchmark against LLMTime [81], a model designed for zero-shot forecasting using LLMs. Following LLMTime, we utilize one sample from each dataset to manage the extensive inference costs. We exclude a related method, Time-LLM [47], from experiments. Time-LLM supports zero-shot learning but requires that the forecasting length and the number of variables/sensors for zero-shot prediction are the same as those used for training.
@@ Appendix H Additional Results: Relation among Prompt Tokens
To compare the difference among tokens before and after training, beyond similarity comparison, we show UMAP plots generated with the prompt tokens before and after training, in Figure 8 and Figure 9. Before training, the prompt tokens from all datasets are dispersed. In contrast, the UMAP of prompt tokens after training reveals that tokens from the same datasets are clustered. However, some tokens from different datasets remain closely positioned, indicating that data from different domains share similar information.
@@ Appendix I Additional Results: Classification Performance Stratified by Datasets
Appendix J Additional Results: Direct Multi-step Forecasting on New Forecasting Lengths
@@
Appendix L Additional Results: Multi-task versus Single-task Learning
To verify the gap between multi-task and single-task learning under fair comparisons, we conduct a experiment to train the single-task models using the same hyper-parameters as the multi-task co-training. As shown in Table 35, multi-task learning achieves stronger performance on both forecasting and classification tasks. Interestingly, under the same hyper-parameters, some classification models fail to converge in the single-task setting, whereas the multi-task model does not have this issue, demonstrating the robustness of multi-task training.
Table 35: Compare UniTS trained by multi-task learning with that trained by single-task learning under same hyper-parameters.
[…]
Multi-task 81.6% 0.439
Single-task 65.3% 0.464
Appendix M Limitations and Future Directions
@@
UniTS primarily aims to unify predictive and generative tasks within a single multi-task model. We demonstrate this by showcasing its adaptability to new data and tasks through prompt learning and few-shot learning. While adapting to new time series data differs fundamentally from generalizing to entirely new data, we will further explore UniTS’s generalization ability for zero-shot learning.
Figure 7: The similarity of prompt tokens among datasets.
Figure 8: UMAP of untrained prompt tokens in UniTS. This plot illustrates that there is no significant organization (clustering) of prompt tokens prior to UniTS training.
Figure 9: UMAP of trained prompt tokens in UniTS. Unlike Figure 8 above, this plot illustrates the meaningful organization (clustering) of prompt tokens by dataset domain category when trained by UniTS.
Appendix N Impact Statement